lindsaylynn420:

Jason had $32. He spent all the money buying four CDs for x dollars each and two magazines for y dollars each. If Jason had bought five CDs and two magazines, he would have run short by $4. The following system of equations models this scenario: 4x + 2y = 32 5x + 2y = 36 Use the system of equations to solve for x and y.

mayathegamergrl:

i would give u the answer- but everyone gots to report me when i do that 😐 so ill try to do it step by step

lindsaylynn420:

okie thanks

mayathegamergrl:

fs

mayathegamergrl:

4x + 2y = 32 5x + 2y = 36 => x = 4 => 4(4) + 2y = 32 => 2y = 16 => y = 8 (is this a good step-by-step?)

thatgirlcrazy:

@mayathegamergrl wrote:
4x + 2y = 32 5x + 2y = 36 => x = 4 => 4(4) + 2y = 32 => 2y = 16 => y = 8 (is this a good step-by-step?)
that's exactly what I was gonna type

lindsaylynn420:

thank you

mayathegamergrl:

@lindsaylynn420 wrote:
thank you
ur welcome (:

KyledaGreat:

To solve for x and y in the system of equations: 4x + 2y = 32 5x + 2y = 36 We can subtract the first equation from the second equation to eliminate y: 5x + 2y - (4x + 2y) = 36 - 32 x = 4 Now that we know x = 4, we can substitute that value into either equation to solve for y. Let's use the first equation: 4(4) + 2y = 32 16 + 2y = 32 2y = 16 y = 8 Therefore, the solution to the system of equations is x = 4 and y = 8. This means that each CD costs $4 and each magazine costs $8.